# Read, Watch and Scream! Sound Generation from Text and Video

### v1

@@ 6 Conclusion
This paper proposes ReWaS, a novel video-and-text-to-sound generation framework. Our main contribution is that audio structural condition, namely energy, is inferred from video to efficiently and effectively input visual condition to the robust T2A model. Therefore, ReWaS can generate complex sounds in the real world without the need for a difficult control design. Quantitative results on VGGSound and Greatest Hits datasets, subjective human study, and qualitative results consistently support that ReWaS can generate natural, temporally well-aligned, and relevant audio for the given video by employing text and video as control.
@@ Appendix A Appendix
A.1 Data preprocessing and feature extraction
@@
During training, we randomly extract 5-second segments from the VGGSound dataset and 2-second segments from the Greatest Hits dataset. However, during the testing phase, we extract video clips ranging from 2 to 7 seconds in duration for the VGGSound dataset, and from 0 to 2 seconds for the Greatest Hits dataset. Video frames are uniformly sampled at 25 fps. Since ReWaS generates audio based on 5-second videos, we duplicated frames from the Greatest Hits dataset to match the length of these 5-second videos. Subsequently, we trimmed the generated audio to a duration of 2 seconds.
@@
We employ SynchFormer [19] trained on VGGSound for the sparse synchronized setting as a video encoder. The video encoder employed in SynchFormer is based on Motionformer [29] pre-trained on Something-Something v2, and fine-tuned on VGGSound and AudioSet. Therefore, the video encoder is strong enough to encode motion dynamics and semantics. We freeze the parameters in the video encoder, and solely train a projection module to estimate energy control. We extract a video feature in the short video clip (0.64 sec). Thus we use a total of 112 length visual embeddings for a 5s video. We note that, for a fair comparison, RGB frames are only used in all methods including ReWaS.
@@
Audios of all videos used in our experiments are resampled to 16kHz sampling rate. We follow the default setting of AudioLDM to compute the mel-spectrogram. Specifically, we use 64-bin mel-spectrograms with 1024 window length. While $f_{\text{min}}$ and $f_{\text{max}}$ are 0 and 8000 respectively, the hop size is 160 and the FFT size is 1024.
To train ReWaS with the onset signal, we follow mostly the same settings as Synchfusion [4]. Specifically, frames are sampled at 15 fps, and resized to 112x112 without image augmentation, resulting in a total of 30 frames per 2 seconds clip. We resample audios to a 16 kHz sampling rate. Using the open-source library Librosa [^1] Module available at https://librosa.org/doc/, we detect the onset signal and interpolate it to 30 frames using nearest-neighbor interpolation to match the frame length.
A.2 Architecture and training details
@@
Energy signal. To encode a video feature into 1-dimensional energy, a projection module $\phi$ consists of a linear layer, two transformer blocks, and MLPs consisting of four FC layers. We use 768 hidden dimensions for the first linear layer and transformer blocks, and the four FC layers’ output dimensions are 128, 64, 16, and 1. The total parameter of $\phi$ is 22M. We choose AudioLDM-M222weights in https://github.com/haoheliu/AudioLDM, and the number of training parameters for fine-tuning AudioLDM with our energy adapter is 182M. ReWaS is optimized by AdamW and the learning rate is fixed to 3e-5 during training. We train ReWaS with 4 V100 GPUs for 33 hours on VGGSound, and 1 hour on GreatestHits respectively.
Onset signal. We re-implement the video onset detection module from Synchfusion [4] as a projection module $\phi$, consisting of a ResNet(2+1)D-18 network followed by a few fully-connected layers. The module generates binary labels using a sigmoid function and applies a threshold of 0.5. Finally, it produces onset signals (0 or 1) for each video frame. To train $\phi$, we adopt the same training settings as Synchfusion. Specifically, we train the video onset detection module for 100 epochs with a batch size of 16, using the AdamW optimizer with a weight decay of $10^{-3}$ and a learning rate of $10^{-4}$. Due to the imbalance of onset sound in the data, we utilize a weighted binary cross-entropy loss [5] with a weight of 10. The onset adapter fine-tuned using AudioLDM-M has the same architecture as the energy adapter, and is trained on Greatest Hits.
A.3 User study
@@
Figure A.1, Figure A.2, and Figure A.3 show the user instructions used in our human evaluation. Before launching Amazon MTurk (AMT), we first conducted an in-lab study with 23 participants; each participant evaluated 20 audio samples for each method and each criterion, namely, they evaluated 240 (20 $\times$ 4 $\times$ 3) generated audio samples. Based on the observation from the in-lab study, we have set the compensation level for each HIT to \$0.45 so that a worker can earn \$15 per hour. At the same time, we observed that a number of participants had trouble keeping focus on the evaluation with 240 samples (each sample takes five seconds). To prevent the low-quality responses from MTurk annotators, we split each evaluation Human Intelligence Task (HIT) on a smaller scale. Each AMT annotator evaluates five audio samples for each method and one additional ground truth audio to prevent random guessing. We published 50 HITs for each criterion, and 150 responses were collected. Finally, we observe that many AMT annotators consistently score high for all questions (e.g., 4 or 5). To ignore noisy responses, we omit responses having an average score larger than 4.0 for 21 questions. 55 responses were omitted after this filtering process.
@@
Figure A.1: User instruction for audio quality (naturalness) test.
@@
Figure A.2: User instruction for video-audio relevance test
@@
Figure A.3: User instruction for temporal alignment test.
A.4 T2A framework
We conduct an ablation study on the T2A model to demonstrate the flexibility of our approach. A key advantage of our method is its modular design, allowing seamless substitution of any T2A diffusion framework. To validate this, we replace the AudioLDM-M [25] backbone with Make-An-Audio [17]. The video encoder used in Make-An-Audio is re-trained to predict the appropriate energy scale of mel-spectogram, which is configured with 80 frequency bins and a hop size of 256 samples, different from the AudioLDM-M configuration. Make-An-Audio is notable for its parameter efficiency, requiring significantly fewer parameters than AudioLDM. This reduction in model complexity translates to substantially shorter training times, with the entire model converging in less than one day. Table A.1 presents a quantitative comparison between the two backbones on VGGSound.
Interestingly, ReWaS built upon Make-An-Audio achieves performance comparable to its AudioLDM-M-based counterpart across most metrics. This underscores the robustness of our framework, demonstrating its ability to maintain high performance even when integrated with a more compact backbone. However, it is worth noting that the Make-An-Audio variant exhibits a slight degradation in overall audio quality such as FID and IS scores. This discrepancy can be attributed to the inherent limitations of the Make-An-Audio architecture, which is computationally efficient but often fails to generate high-quality audio over large-scale T2A frameworks. In other words, the development of T2A models could advance our model.
Table A.1: Performance comparison on VGGSound [2] with different T2A framework.
@@ A.4 T2A framework
ReWaS (Make-An-Audio) 34.55 6.35 4.76 42.37 0.75 74.3M
ReWaS (AudioLDM) 29.09 15.96 6.37 41.70 1.29 204M
A.5 More qualitative results
We illustrate estimated energy from video in Figure A.4. The results show the correlation between our energy control generated from video and GT energy obtained from reference audio.
We provide more examples to show the effectiveness of ReWaS qualitatively. Figure A.5 is another qualitative example that demonstrates the ReWaS can generate multiple semantics without providing the context in textual prompts. In Figure A.6, ReWaS successfully generates "silent" audio sounds with visual input. In contrast, other baseline models fail to produce the corresponding sound (e.g., alarm clock ringing) due to misaligned visual and sound contexts, often generating unintended sounds or remaining silent when they should produce sound. This suggests that baseline models may resort to generating random sounds when faced with a change in frame scene since the visual context and sound context are misaligned but they also generate the sound where they have to be silent. Figure A.7 shows the effectiveness of ReWaS in aligning sound with corresponding frames, particularly in scenarios where man-made video-audio data lacks correlation due to the sound being added later to the frames. ReWaS exhibits superior performance in generating suitable audio sounds with better temporal synchronization accurately capturing small movements of objects (e.g., lip synchronization). For example, the energy of the growling sound becomes bigger as the lion opens its mouth. Lastly, Figure A.8 illustrates ReWaS’s ability to generate temporally synchronized sounds, further highlighting its effectiveness.
Figure A.4: Examples of energy controls from input videos.
Figure A.5: Additional example of demonstrating the effectiveness of video input. ReWaS can make the sound missing in the text prompt (‘spray’), while simultaneously aligning multiple sounds with various semantics to the video.
Figure A.6: Example of audio sound from misaligned visual input. ReWaS can make the desired sound and the silent moment like ground-truth sound.
Figure A.7: Example of audio with improved synchronization, capturing small movements (e.g., a lion’s lip synchronization).
Figure A.8: Examples of generated audio sounds demonstrating the capability of temporal synchronization.

### v2

@@ 6 Conclusion
This paper proposes ReWaS, a novel video-and-text-to-audio generation framework. Our key idea lies in inferring audio structural condition, namely energy, from video to efficiently and effectively input the visual condition to the robust T2A model. Therefore, ReWaS can generate complex sounds in the real world without the need for a difficult control design. Quantitative results on VGGSound and Greatest Hits datasets, subjective human study, and qualitative results consistently support that ReWaS can generate natural, temporally well-aligned, and relevant audio for the given video by employing text and video as control.
Acknowledgements
The NAVER Smart Machine Learning (NSML) platform (Kim et al. 2018) had been used for experiments.
@@ Appendix A Appendix
A.1 Data preprocessing
@@
During training, we randomly extract 5-second segments from the VGGSound dataset (Chen et al. 2020) and 2-second segments from the Greatest Hits dataset. However, during the testing phase, we extract video clips ranging from 2 to 7 seconds in duration for the VGGSound dataset, and from 0 to 2 seconds for the Greatest Hits dataset (Owens et al. 2016). Video frames are uniformly sampled at 25 fps. Since ReWaS generates audio based on 5-second videos, we duplicated frames from the Greatest Hits dataset to match the length of these 5-second videos. Subsequently, we trimmed the generated audio to a duration of 2 seconds.
For comparison with baselines, SpecVQGAN (Iashin and Rahtu 2021), Diff-Foley (Luo et al. 2024) and Seeing&Hearing (Xing et al. 2024) (10s, 8s, and 10s, respectively) for the test videos. Then, we extract the 5-second clip corresponding to the same video frames used in our method. Since Im2wav (Sheffer and Adi 2023) is designed to generate sound with a fixed length of 4 seconds, we first generate the initial 4 seconds and extend it by generating an additional 1 second, resulting in a 5-second audio clip.
A.2 Feature extraction
@@
Video features. We employ SynchFormer (Iashin et al. 2024) trained on VGGSound  (Chen et al. 2020) for the sparse synchronized setting as a video encoder. The video encoder employed in SynchFormer is based on Motionformer (Patrick et al. 2021) pre-trained on Something-Something v2 (Goyal et al. 2017), and fine-tuned on VGGSound and AudioSet (Gemmeke et al. 2017). Therefore, the video encoder is strong enough to encode motion dynamics and semantics. We freeze the parameters in the video encoder, and solely train a projection module to estimate energy control. We extract a video feature in the short video clip (0.64 sec). Thus we use a total of 112 length visual embeddings for a 5s video. We note that, for a fair comparison, RGB frames are only used in all methods including ReWaS.
@@
Audio features. Audios of all videos used in our experiments are resampled to 16kHz sampling rate. We follow the default setting of AudioLDM to compute the mel-spectrogram. Specifically, we use 64-bin mel-spectrograms with 1024 window length. While $f_{\text{min}}$ and $f_{\text{max}}$ are 0 and 8000 respectively, the hop size is 160 and the FFT size is 1024.
A.3 Architecture and training details
Test dataset. For our experiments, we leverage a subset of 160k videos from VGGSound (Chen et al. 2020) due to the availability of public videos at the time of training. We split the train data list into training and validation subsets following SpecVQGAN (Iashin and Rahtu 2021).
@@ A.3 Architecture and training details
Energy signal. To encode a video feature into 1-dimensional energy, a projection module $\phi$ consists of a linear layer, two transformer blocks, and MLPs consisting of four FC layers. We use 768 hidden dimensions for the first linear layer and transformer blocks, and the four FC layers’ output dimensions are 128, 64, 16, and 1. The total parameter of $\phi$ is 22M. We choose AudioLDM-Mweights in https://github.com/haoheliu/AudioLDM, and the number of training parameters for fine-tuning AudioLDM (Liu et al. 2023) with our energy adapter is 182M. ReWaS is optimized by AdamW and the learning rate is fixed to 3e-5 during training. We train ReWaS with 4 V100 GPUs for 33 hours on VGGSound, and 1 hour on Greatest Hits (Owens et al. 2016) respectively.
Details of Make-An-Audio backbone framework. The video encoder used in Make-An-Audio (Huang et al. 2023b) is re-trained to predict the appropriate energy scale of mel-spectogram, which is configured with 80 frequency bins and a hop size of 256 samples, different from the AudioLDM-M configuration. Make-An-Audio is notable for its parameter efficiency, requiring significantly fewer parameters than AudioLDM. This reduction in model complexity translates to substantially shorter training times, with the entire model converging in less than one day.
A.4 User study
@@
Figure 8, Figure 9, and Figure 10 show the user instructions used in our human evaluation. Before launching Amazon MTurk (AMT), we first conducted an in-lab study with 23 participants; each participant evaluated 20 audio samples for each method and each criterion, namely, they evaluated 240 (20 $\times$ 4 $\times$ 3) generated audio samples. Based on the observation from the in-lab study, we have set the compensation level for each HIT to \$0.45 so that a worker can earn \$15 per hour. At the same time, we observed that a number of participants had trouble keeping focus on the evaluation with 240 samples (each sample takes five seconds). To prevent the low-quality responses from MTurk annotators, we split each evaluation Human Intelligence Task (HIT) on a smaller scale. Each AMT annotator evaluates five audio samples for each method and one additional ground truth audio to prevent random guessing. We published 50 HITs for each criterion, and 150 responses were collected. Finally, we observe that many AMT annotators consistently score high for all questions (e.g., 4 or 5). To ignore noisy responses, we omit responses having an average score larger than 4.0 for 21 questions. 55 responses were omitted after this filtering process.
@@
Figure 8: User instruction for audio quality (naturalness) test.
@@
Figure 9: User instruction for video-audio relevance test
@@
Figure 10: User instruction for temporal alignment test.
A.5 More qualitative results
Energy controls from videos. We illustrate estimated energy from video in Figure 11. The results show the correlation between our energy control generated from video and GT energy obtained from reference audio.
Effectiveness of the text prompt. In the input video shown in Fig. 12, rain streaks are barely visible, while we want to hear the sound of rain. ReWaS can emphasize the desired sound with the help of a text prompt ‘raining’. While V2A methods struggle to generate corresponding sound, Seeing&Hearing (Xing et al. 2024) and ReWaS can effectively calibrate the semantics by user text prompt. However, ReWaS shows better CLAP score (Wu* et al. 2023) than Seeing&Hearing. Furthermore, we assert that Seeing&Hearing is heavily dependent only on text prompt. For example, as shown in Figure 13, if there are redundant frames, ReWaS can only successfully calibrate the semantics with textual prompt but also it generates “silent” audio sounds when there is a scene change. In contrast, other baseline models such as SpecVQGAN (Iashin and Rahtu 2021), Im2wav (Sheffer and Adi 2023) and Diff-Foley (Luo et al. 2024) fail to produce the corresponding sound (e.g., alarm clock ringing) due to misaligned visual and sound contexts, often generating unintended sounds or remaining silent when they should produce sound. Although Seeing&Hearing (Xing et al. 2024) can produce corresponding sounds, it fails to generate “silent” audio when there is a change in visual scenes. This suggests that baseline models may either resort to generating random sounds when faced with a scene change due to misaligned visual and sound contexts, or they produce sound when they should remain silent, ignoring the visual context.
Effectiveness of visual control. Figure 14 and Figure 15 are examples when energy signal serving additional temporal information. As shown in  Figure 14, when a person talking and playing a dart game in an input video, the original AudioLDM (Liu et al. 2023) generates only the sound of talking, ignoring ‘dart’ prompt. Additionally, aligning generated sound with video is challenging in AudioLDM. In comparison, ReWaS not only generates both the sound of talking and dart but also aligns the sound with the frames. Figure 15 presents another example. Unlike AudioLDM, which repeatedly generates the same spray and car engine sounds, ReWaS accurately captures the spray sound at the right moment thanks to the visual control without additional text prompt ‘spray’. Furthermore, the result demonstrates the limitation of T2A methods for automatic Foley synthesis, because they cannot watch a video. This demonstrates the effectiveness of visual control by ReWaS.
Subtle visual movements. Figure 16 and Figure 17 demonstrate the effectiveness of ReWaS in aligning sound with corresponding frames, achieving temporal alignment by accurately capturing small object movements, such as lip synchronization. As illustrated in Figure 16, the the intensity of growling sound increases as the lion opens its mouth. In another example of Figure 17, ReWaS also produces temporally synchronized sounds with mouth movements, underscoring its overall effectiveness.
General text prompt. Figure 18 provides an example that evaluates the capability of ReWaS using more general text prompts. We generate audio samples with another generated video from KLINGhttps://kling.kuaishou.com/en. Our method is the only one that captures the increasing intensity of the sound as the onions are cut from the edge to the center. Both the T2A model, AudioLDM, and the V2T&V2A model, Seeing&Hearing, can generate corresponding sounds, but they lack visual temporal alignment in the generated results.
A.6 Extended analysis
Isolating and editing energy across different events. We observe that visual controls provide temporal information, raising the following question of whether it is possible to isolate the energy associated with each event or edit the sound on a single track. To address this, we combine two videos with distinctly different styles within the same category. As shown in Figure 19, we observe that the generated sounds change their rhythm or tempo accordingly and average energy MAE across events shows 2-3 times difference, indicating that isolating or editing energy by event is feasible.
Roles of text prompts and visual energy control. In Table 6, we present additional experimental results on subset of VGGSound (Chen et al. 2020) to explore the specific roles of text prompts and visual energy control. The first and second columns show results using null text and misaligned text, respectively, while the third column shows results without visual energy control. For misaligned text prompts, we randomly shuffle text prompts from different video categories. We observe that the model typically generates the main audio sound based on the text prompt while aligning the sound’s temporal dynamics with the video movement. In particular, misaligned or null text prompts result in a much lower CLAP score (Huang et al. 2023b), while the AV-align (Yariv et al. 2024) score shows only a slight decrease from 0.30 to 0.27, highlighting the importance of accurate text guidance. In contrast, omitting energy control results in significant drops in the AV-align score, demonstrating that energy control ensures temporal alignment between generated audio and video. In conclusion, the text and energy controls each play critical and distinct roles in generating semantically and temporally aligned audio outputs.
Table 6: Ablation Studies on text prompt and energy controls.
Metrics Ours w/o Text misaligned Text w/o Energy
@@ A.6 Extended analysis
[…]
A.7 Future research directions
Our architecture is flexible and can also adapt to TTS models, like Diff-TTS (Jeong et al. 2021). Our model could potentially generate nuanced emotional intonation by facial movements, which could have a powerful impact given the limitations in emotional video-to-speech generation. Music generation from dance videos is also promising future research direction.
A.8 Limitations
Although our approach successfully leverages the text and video control simultaneously, our method shares the limitation of AudioLDM, namely, hallucination in generated samples. For example, for a given “basketball bounce” video, ReWaS generates a squeaking sound, even if the player is standing still. We conduct a simple human evaluation on 100 random generated audio samples by ReWaS on a 5-point scale (1 = no hall., 5 = significant hall.). We got 2.1 average score, with hallucinations identified in 18% of cases. This problem might be mitigated if we can use a better AudioLDM model.
Figure 11: Examples of energy controls from input videos.
Figure 12: Effectiveness of text prompt. Videos in the real world are sometimes noisy. For example, when videos are hard to distinguish the semantics, text prompts used in ReWaS calibrate the results. Text prompt is “raining”.
Figure 13: Example of audio sound from misaligned visual input. ReWaS can make the desired sound and make the silent moment like ground-truth sound.
Figure 14: Effectiveness of video input. In ReWaS, energy control from video input transfers additional temporal information.
Figure 15: Additional example of effectiveness of video input.
Figure 16: Example of audio with improved synchronization, capturing small movements (e.g., a lion’s lip synchronization).
Figure 17: Examples of generated audio sounds demonstrating the capability of temporal synchronization.
Figure 18: Example of general user prompt.
Figure 19: Example of concatenation of different video event.